OpenStudy (anonymous):

What is the slope of the graph of 2x – 6y = 9? A.3 B.-3 C.-1/3 D.1/3

OpenStudy (anonymous):

2x-6y = 9 Rearranging: 6y = 2x - 9 Divide through by 6: y = (1/3)x - 3/2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Thank y'all

OpenStudy (gudden):

@LogicFail and @kamdenyoung ... You guys look new to the site... Please make sure to explain everything regarding the concept, so that the one who asked the question gets to know the actual concept and can do other such questions,, After all, It's a study site :)

OpenStudy (anonymous):

First you have to change the equation in the form of 'y = mx + b'. 2x – 6y = 9 -6y = -2x + 9 y = 0.333333x + 1.5 y = mx + b Where 'm' = slope. So what is the slope @destinistmyers ?

OpenStudy (gudden):

@destinistmyers ..... Slope = y axis / x axis... So next time, just re-arrange these variables in the ratio of y axis to that of x axis..:) And you would the slope...

OpenStudy (anonymous):

0.3333333 is the same thing as 1/3! So your answer is D.

OpenStudy (anonymous):

Ok, I just Thought You Told Them The Answer With Working.

OpenStudy (anonymous):

Thank you guys

OpenStudy (anonymous):

I Said D In My Working 2x-6y = 9 Rearranging: 6y = 2x - 9 Divide through by 6: y = (1/3)x - 3/2 y= (1/3)

OpenStudy (anonymous):

Not A Problem.

OpenStudy (gudden):

@LogicFail .. that's correct that you can tell them the answer by working.... but the condition is it that, if the user isn't able to get...what you say...:) Otherwise... I appreciate your spirit,..! Just make sure next time..:)

OpenStudy (anonymous):

Will Do Gudden.

OpenStudy (gudden):

:)

OpenStudy (anonymous):

I just needed help because i i had 70% on this homework if i don't get above and 80% on every piece of homework it doesn't count and i have to redo it :/

OpenStudy (gudden):

Should I explain it to you again ?

OpenStudy (anonymous):

No i meant i had a 70% on it that's why I need help to get it higher

OpenStudy (gudden):

Anyway... I gotta go... Good Night...

OpenStudy (anonymous):

night